OpenStudy (icalibear):

what is the side length of a square with an area of 144x^2? How do i set up that equation?

OpenStudy (anonymous):

since the area is \(\large 144x^2 \), you'll need to take the square root of that expression to get the length of the side of the square: side = \(\large \sqrt{144x^2} \)

OpenStudy (icalibear):

12?

OpenStudy (anonymous):

formula for area of square is Area = side ^2 => 144x^2 = side square now solve by taking squareroot of 144x^2

OpenStudy (icalibear):

so its 12 x

OpenStudy (icalibear):

is it 12x or just 12?

OpenStudy (anonymous):

correct...☺

OpenStudy (anonymous):

now check: will (12x)(12x) = 144x^2 ?? if yes, then you're right....

OpenStudy (icalibear):

yep c:

OpenStudy (anonymous):

it is 12x